# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ Abstract
We study simultaneous symmetry-breaking in a spontaneous Floquet state, focusing on the specific case of an atomic condensate. We first describe the quantization of the Nambu-Goldstone (NG) modes for a stationary state simultaneously breaking several symmetries of the Hamiltonian by invoking the generalized Gibbs ensemble, which enables a thermodynamical description of the problem. The quantization procedure involves a Berry-Gibbs connection, which depends on the macroscopic conserved charges associated to each broken symmetry and whose curvature is not invariant under generalized gauge transformations. We extend the formalism to Floquet states simultaneously breaking several symmetries, where Goldstone theorem translates into the emergence of Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In the case of a spontaneous Floquet state, there is a genuine temporal FNG mode arising from the continuous time-translation symmetry breaking, whose quantum amplitude provides a rare realization of a time operator in Quantum Mechanics. Furthermore, since they conserve energy, spontaneous Floquet states can be shown to possess a conserved Floquet charge. Both the temporal FNG mode and the Floquet charge are distinctive features of a spontaneous Floquet state, absent in conventional, driven systems. Nevertheless, these also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. We apply our formalism to a particular realization of spontaneous Floquet state, the CES state, which breaks $U(1)$ and time-translation symmetries, representing a time supersolid. Using the Truncated Wigner method, we numerically compute its quantum fluctuations, which are theoretically predicted to be dominated by the temporal FNG mode at long times, observing a remarkable agreement between simulation and theory. Based on these results, we propose a feasible experimental scheme to observe the temporal FNG mode of the CES state.
I Introduction
@@
Noether and Goldstone theorems are two of the most fundamental results in Physics, and can be regarded as the sides of the same coin Callen (1985). The former states that any continuous symmetry is translated into a conservation law, while the latter asserts that solutions spontaneously breaking one of those symmetries have associated a zero-energy mode, known as Nambu-Goldstone (NG) mode. Spontaneous symmetry breaking is still a hot topic of research, motivating the search for novel phases of quantum matter which simultaneously break several symmetries, such as supersolids Léonard et al. (2017); Li et al. (2017), quantum Hall ferromagnets Ezawa (2008); Kharitonov (2012); Kim et al. (2021); Coissard et al. (2022), spinor Stenger et al. (1998); Stamper-Kurn and Ueda (2013); Blinova et al. (2023) and rotating Polkinghorne and Simula (2021) condensates, or quantum droplets Chomaz et al. (2016); Cabrera et al. (2018); Liu et al. (2019); Zin et al. (2021).
@@
A fascinating perspective is provided by time crystals Wilczek (2012); Sacha (2020), which spontaneously break time-translation symmetry, perhaps the most fundamental symmetry in nature. Although a no-go theorem ruling out them as first conceived was proven Watanabe and Oshikawa (2015), time crystals are still possible in the out-of-equilibrium arena. Discrete Choi et al. (2017); Zhang et al. (2017); Smits et al. (2018); Rovny et al. (2018); Kyprianidis et al. (2021); Randall et al. (2021); Mi et al. (2022); Frey and Rachel (2022) and continuous time crystals Autti et al. (2018); Kongkhambut et al. (2022); Liu et al. (2023); Greilich et al. (2024) have been observed in a wide variety of systems, ranging from cold atoms to superconducting quantum computers. In particular, discrete time crystals emerge in Floquet systems as a subharmonic response to the external periodic driving Sacha (2015); Else et al. (2016). In general, Floquet systems Shirley (1965); Sambe (1973); Grifoni and Hänggi (1998) provide rich scenarios to study a number of out-of-equilibrium features such as prethermalization Peng et al. (2021), topological insulation Lindner et al. (2011), dynamical phase transitions Prosen and Ilievski (2011), high-harmonic generation Murakami et al. (2018), or protected cat states Pieplow et al. (2019) and flat-band superfluidity Mateos et al. (2023).
@@
It was recently suggested that Floquet physics can also arise in time-independent configurations, where the system self-consistently oscillates as a Floquet state of its effective Hamiltonian due to many-body interactions de Nova and Sols (2022). A specific realization of spontaneous Floquet state was proposed using an atomic Bose-Einstein condensate (BEC), the so-called continuous emission of solitons (CES) state de Nova et al. (2016, 2021). As a result, the CES state breaks $U(1)$ and continuous time-translation symmetries, providing the temporal analogue of a supersolid, which has been labeled as time supersolid Autti et al. (2018). Some natural questions then arise: how to combine several broken symmetries when one of them is time itself? What is the temporal version of Goldstone theorem?
@@
In this work, we study simultaneous symmetry breaking in spontaneous Floquet states, focusing on the specific case of an atomic BEC close to zero temperature for illustrative purposes. We first discuss simultaneous symmetry breaking in stationary states within the framework of the generalized Gibbs ensemble Cazalilla (2006); Rigol et al. (2007); Langen et al. (2015), which emerges in the study of thermalization in isolated systems, where each conserved charge further restricts the available phase-space for the dynamics. The equitative treatment of each broken symmetry provided by the generalized Gibbs ensemble allows for a general and complete description of the problem, including its thermodynamics.
@@
The NG modes emerge as zero-energy modes of the spectrum of collective excitations. Each NG mode is paired with a Gibbs mode, which accounts for the fluctuations of the quantum state with respect to the associated conserved charge. These fluctuations stem from the fact that a symmetry-broken state cannot be an eigenstate of the charge generating the corresponding symmetry transformation. Remarkably, these Goldstone-Gibbs modes can be elegantly described by a geometrical approach that involves a generalization of the Berry connection defined within a manifold whose variables are the conserved charges, denoted as a result as the Berry-Gibbs connection. This connection is extended to the full manifold which includes the continuous parameters of the broken symmetries, where its curvature gives rise to a symplectic form that provides the commutation relations between the amplitudes of the Goldstone-Gibbs modes, hence behaving as conjugate coordinate-momenta. This relation is also translated to the dynamics, where the amplitude of the Gibbs modes behaves as a conserved momentum, since it represents the quantum fluctuations of the corresponding charge, while the amplitude of the Goldstone modes displays a ballistic motion, with a velocity proportional to the momenta. Moreover, by invoking only thermodynamic considerations, we show that the Goldstone-Gibbs sector can be diagonalized in terms of a set of uncoupled NG modes, which determine how the original NG modes hybridize within the different branches of the spectrum in the low-frequency limit. Our results for simultaneous symmetry breaking in a stationary state thus generalize the original works of Refs. Lewenstein and You (1996); Dziarmaga (2004) to an arbitrary number of broken symmetries, and provide a deeper understanding of the nature of the Goldstone-Gibbs modes.
@@
As an example, we apply our formalism to a cnoidal wave in a superfluid, a model that has been used to understand fundamental aspects of supersolidity Martone et al. (2021), validating our theoretical results. An important conclusion of this study is that periodic motion along a closed orbit spawned by the broken generators may result in the misidentification of time-crystalline behavior, not representing a genuine symmetry-breaking of time-translation invariance.
@@
The generality of the developed framework allows its application for the study of simultaneous symmetry breaking in Floquet states, where the NG modes are now translated into Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In particular, since spontaneous Floquet states break continuous time-translational symmetry, they present a genuine temporal FNG mode. For a complete characterization of the problem, we develop the $(t,\phi)$ formalism, denoted in this way because of its analogy with the conventional $(t,t^{\prime})$ formalism Peskin and Moiseyev (1993); Heimsoth et al. (2012). In combination with the generalized Gibbs ensemble, the $(t,\phi)$ formalism leads to a thermodynamic description analogous to that of stationary states, which we term as Floquet thermodynamics. For spontaneous Floquet states, this is a direct consequence of energy conservation, which implies the existence of a conserved Floquet charge. Conventional Floquet states also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. This is because driven systems are isoperiodic, i.e., they operate at an externally fixed frequency, which is the conjugate variable of the Floquet charge, in analogy with isobaric and isothermal systems where the environment fixes the pressure or the temperature, respectively. The temporal FNG mode and the Floquet charge are thus the distinctive features of a spontaneous Floquet state, absent in driven systems.
@@
The quantization of the FNG and Gibbs modes is performed along the same lines as in the stationary case, described by a Berry-Gibbs connection that results in a similar coordinate-momentum correspondence. However, a novel feature is that the amplitude of the temporal FNG mode can be regarded as an effective time coordinate, whose conjugate momentum represents the energy fluctuations. Therefore, the quantum amplitude of the temporal FNG mode represents a unique realization of a time operator in Quantum Mechanics.
@@
Finally, we apply our formalism to the CES state, computing its quantum fluctuations with the help of the Truncated Wigner method Sinatra et al. (2002); Carusotto et al. (2008). Theory predicts that, at long times, the temporal FNG mode dominates the density-density correlations, observing an excellent agreement with the numerical results. Based on these results, we propose a feasible experimental scheme to observe the temporal FNG mode of the CES state.
@@
The article is arranged as follows. Section II discusses how the general results from Goldstone theorem are recovered in variational approaches in many-body systems. Section III addresses simultaneous symmetry breaking in stationary states within the generalized Gibbs ensemble. Section IV applies the formalism of Sec. III to a cnoidal wave in a superfluid. Section V extends the results of Sec. II to Floquet states, both spontaneous and conventional. Section VI discusses simultaneous symmetry breaking in Floquet systems. Section VII applies the formalism of Sec. VI to study the quantum fluctuations of the CES state. Technical details are presented in the Appendix.
II Variational Goldstone Theorem
@@
Originally derived within the framework of relativistic quantum field theory Goldstone et al. (1962), the general idea behind Goldstone theorem is universal, namely, solutions which break some continuous symmetry of the equations of motion include a linear zero mode associated to the broken symmetry. This feature is even present in classical mechanics. Consider a particle of mass $m$ moving within a potential $V(\mathbf{x})$ which is invariant under some continuous symmetry transformation of the coordinates, $V(\mathbf{x}^{\prime})=V(\mathbf{x})$, $\mathbf{x}^{\prime}=\mathbf{x}^{\prime}(\mathbf{x},\alpha)$, where $\alpha$ parameterizes the transformation in such a way that $\alpha=0$ corresponds to the identity transformation, $\mathbf{x}^{\prime}(\mathbf{x},0)=\mathbf{x}$. As standard in Lie theory, the symmetry transformation can be expanded as
@@
where the vector $\boldsymbol{\xi}=-iL\mathbf{x}$ describes the symmetry transformation at the infinitesimal level and $L$ is the generator of the continuous transformation on the coordinates. At the same time, Noether’s theorem guarantees the existence of a conserved quantity associated to that symmetry.
Time-independent solutions to the equations of motion are given by the equilibrium points $\mathbf{x}_{0}$ which extremize the potential, $\boldsymbol{\nabla}V(\mathbf{x}_{0})=0$. Linearized motion around equilibrium positions is characterized by the eigenvalues $m\Omega^{2}_{i}$ of the matrix $\partial^{2}_{ij}V(\mathbf{x}_{0})$. In particular, potential minima are stable equilibrium points, as then all $\Omega_{i}$ are real and represent the frequencies of the normal modes of oscillation.
@@
$\partial^{2}_{ij}V(\mathbf{x}_{0})\xi_{0}^{j}=0,~{}\boldsymbol{\xi}_{0}=-iL\mathbf{x}_{0},$ (2)
@@
where hereafter we use Einstein convention summation unless otherwise stated. Thus, if the symmetry is broken by the equilibrium position, which means that is not invariant under the symmetry transformation, $L\mathbf{x}_{0}\neq 0$, then there is a zero-frequency normal mode, the Nambu-Goldstone mode. Qualitatively, this zero mode results from the fact that there is no restoring force along the orbit of the equilibrium point under the symmetry transformation because potential energy is conserved there.
@@
A simple example to visualize this behavior is given by a particle in a 2D $y$-independent potential, $V(x,y)=V(x)$, with $V(0)$ a global minimum and $V^{\prime\prime}(0)=m\omega_{x}^{2}$. In that case, if we consider that the particle is placed at the equilibrium position $\mathbf{x}=(0,y_{0})$, small displacements along the $x$-axis lead to oscillations with frequency $\omega_{x}$. However, due to the symmetry of the problem, $\omega_{y}=0$, and any small displacement along the $y$-axis leads to an unbounded motion with constant velocity, $y(t)=y_{0}+v_{y}t$, since there is no restoring force in that direction.
The previous concepts can be straightforwardly extended to quantum many-body systems. In that context, it is common to use variational approaches based on trial wave functions characterized by a reduced number of parameters in order to tackle the exponentially large complexity of the problem. For example, in the Dirac-Frenkel variational principle, the ansatz state |$\ket{\Psi(t)}$ is required to extremize
$L(t)\equiv\bra{\Psi(t)}i\hbar\partial_{t}-\hat{H}\ket{\Psi(t)},$ (3)
which would yield the exact Schrödinger equation if no restriction were imposed on |$\ket{\Psi(t)}$  [^1] Technically, the extremization condition is $\bra{\delta\Psi(t)}i\hbar\partial_{t}-\hat{H}\ket{\Psi(t)}=0$. The Dirac-Frenkel formalism can also be extended to finite temperatures DeAngelis and Gatoff (1991). Other variational methods are based on a Lagrangian approach, where the trial wave function extremizes some action; indeed, Eq. (3) can be regarded as an effective Lagrangian. A large number of well-known equations fall under this class of approximation, including the Gross-Pitaevskii (GP) equation Pitaevskii and Stringari (2016), the Hartree-Fock (HF) equations Giuliani and Vignale (2005), the Gutzwiller ansatz Jaksch et al. (1998), or the MultiConfiguration Time-Dependent Hartree (MCTDH) method Caillat et al. (2005); Alon et al. (2008). In all cases, if we group the variational parameters in a generic vector $X$, the resulting equations of motion take the form
$i\hbar\frac{dX^{i}}{dt}=H^{i}_{j}(X)X^{j},$ (4)
@@
where the vector components $X^{i}$ can be orbital wave functions (GP and HF equations), wave-function coefficients (Gutzwiller ansatz), or even both (MCTDH method). The matrix $H^{i}_{j}(X)$ can be nonlinear in the vector components $X^{i}$, acting as the effective Hamiltonian of the problem. For the present moment, we take the original Hamiltonian $\hat{H}$ as time-independent, so $H^{i}_{j}(X)$ only depends implicitly on time through $X$.
@@
Now, we assume that there are stationary solutions of the form $X^{i}(t)=X^{i}e^{-i\epsilon^{i}t/\hbar}$, satisfying the self-consistent eigenvalue equation
$H^{i}_{j}(X)X^{j}=\epsilon^{i}X^{i}.$ (5)
@@
The spectrum of collective modes associated to this stationary solution is obtained by considering small perturbations $X^{i}(t)=X^{i}+\delta X^{i}(t)$, which yields a linear equation of the general form
$i\hbar\frac{d\delta X^{i}}{dt}=M^{i}_{j}(X)\delta X^{j},$ (6)
@@
with $M^{i}_{j}(X)$ some matrix that depends on the background stationary solution. The eigenvalues of $M^{i}_{j}(X)$ are the energies of the collective modes. This is indeed the case of the Bogoliubov-de Gennes (BdG) equations for both the GP equation Pitaevskii and Stringari (2016) and the Gutzwiller ansatz Caleffi et al. (2020), and of the Time-Dependent Hartree-Fock Approximation (TDHFA) for the HF equations Thouless (2014).
@@
In the same fashion as in the classical example, if the stationary solution $X$ spontaneously breaks a continuous symmetry of the problem, so $X^{\prime}=X^{\prime}(X,\alpha)=X-i\alpha LX+\ldots$ is also a stationary solution, then the vector $LX\neq 0$ is a zero mode of the linearized equations of motion,
$0=M^{i}_{j}(X)(LX)^{j}.$ (7)
@@
This is the NG mode associated to the spontaneous symmetry breaking. We note that this derivation also applies to pseudo-NG modes Weinberg (1972); Georgi and Pais (1975), which emerge from extended symmetries in the equations of motion that were not present in the original Hamiltonian.
III Simultaneous symmetry breaking in many-body field theory

### v4

# Simultaneous symmetry breaking in spontaneous Floquet states: temporal Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

@@ Abstract
We study simultaneous symmetry breaking in spontaneous Floquet states, focusing on the specific case of an atomic condensate. We first describe the quantization of the Nambu-Goldstone (NG) modes for a stationary state simultaneously breaking several symmetries of the Hamiltonian by invoking the generalized Gibbs ensemble, which enables a thermodynamical description of the problem. The quantization procedure involves a Berry-Gibbs connection, which depends on the macroscopic conserved charges associated to each broken symmetry and whose curvature is not invariant under generalized gauge transformations. We extend the formalism to Floquet states, where Goldstone theorem translates into the emergence of Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In the case of a spontaneous Floquet state, there is a genuine temporal FNG mode arising from the continuous time-translation symmetry breaking, whose quantum amplitude provides a rare realization of a time operator in Quantum Mechanics. Furthermore, since they conserve energy, spontaneous Floquet states can be shown to possess a conserved Floquet charge. Both the temporal FNG mode and the Floquet charge are distinctive features of spontaneous Floquet states, absent in conventional, driven systems. Nevertheless, as these operate at fixed frequency, they also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. An important conclusion of our analysis is that certain time crystals may not represent a genuine spontaneous temporal symmetry-breaking, since their periodic behavior just results from constant motion along a closed orbit spawned by the generators of other broken symmetries. We apply our formalism to a particular realization of spontaneous Floquet state, the so-called state of continuous emission of solitons (CES), which breaks $U(1)$ and time-translation symmetries, representing a time supersolid. We numerically compute its density-density correlations, predicted to be dominated by the temporal FNG mode at long times, observing a remarkable agreement between simulation and theory. Based on these results, we propose a feasible experimental scheme to observe the temporal FNG mode of the CES state.
1 Introduction
@@
Noether and Goldstone theorems are two of the most fundamental results in Physics, and can be regarded as the sides of the same coin [1]. The former states that any continuous symmetry is translated into a conservation law, while the latter asserts that solutions spontaneously breaking one of those symmetries have associated a zero-energy mode, known as Nambu-Goldstone (NG) mode. Spontaneous symmetry breaking is still a hot topic of research, motivating the search for novel phases of quantum matter which simultaneously break several continuous symmetries, such as supersolids [2, 3], quantum Hall ferromagnets [4, 5, 6, 7], spinor [8, 9, 10] and rotating [11] condensates, or quantum droplets [12, 13, 14, 15].
@@
A fascinating perspective is provided by time crystals [16, 17], which spontaneously break time-translation symmetry, perhaps the most fundamental symmetry in nature. Although a no-go theorem ruling out them as first conceived was proven [18], time crystals are still possible in the out-of-equilibrium arena. Discrete [19, 20, 21, 22, 23, 24, 25, 26] and continuous time crystals [27, 28, 29, 30, 31] have been observed in a wide variety of systems, ranging from cold atoms to superconducting quantum computers. In particular, discrete time crystals emerge in Floquet systems as a subharmonic response to the external periodic driving [32, 33]. In general, Floquet systems [34, 35, 36] provide rich scenarios to study a number of out-of-equilibrium features such as prethermalization [37], topological insulation [38], dynamical phase transitions [39], high-harmonic generation [40], or protected cat states [41] and flat-band superfluidity [42].
@@
It was recently shown that Floquet physics can also arise in time-independent configurations, where the system self-consistently oscillates as a Floquet state of its effective Hamiltonian due to many-body interactions [43]. A specific realization of spontaneous Floquet state was proposed using an atomic Bose-Einstein condensate (BEC), the so-called state of continuous emission of solitons (CES) [44, 45]. As a result, the CES state breaks $U(1)$ and continuous time-translation symmetries, providing the temporal analogue of a supersolid, which has been labeled as time supersolid [27]. Some natural questions then arise: how to combine several broken symmetries when one of them is time itself? What is the temporal version of Goldstone theorem?
@@
In this work, we study simultaneous symmetry breaking (here defined as the spontaneous symmetry breaking of several continuous symmetries by the same state) in spontaneous Floquet states, focusing on the specific case of an atomic BEC close to zero temperature for illustrative purposes. We first discuss simultaneous symmetry breaking in stationary states within the framework of the generalized Gibbs ensemble [46, 47, 48], since its equitable treatment of each broken symmetry allows for a general and complete description of the problem, including its thermodynamics. The generalized Gibbs ensemble emerges in the study of thermalization in isolated systems, where each conserved charge further restricts the available phase-space for the dynamics, leading to a violation of the eigenstate thermalization hypothesis [49, 50, 51], in parallel to other mechanisms such as many-body localization [52, 53] or Hilbert space fragmentation [54].
@@
The NG modes appear as zero-energy modes of the spectrum of collective excitations. Each NG mode is paired with a Gibbs mode, which accounts for the fluctuations of the quantum state with respect to the associated conserved charge, stemming from the fact that a symmetry-broken state cannot be an eigenstate of the charge generating the corresponding symmetry transformation. Remarkably, the Goldstone-Gibbs modes can be described by a geometrical approach that involves a generalization of the Berry connection, defined within a manifold whose variables are the conserved charges, denoted as a result as the Berry-Gibbs connection. This connection is extended to the full manifold which includes the continuous parameters of the broken symmetry transformations, where its curvature gives rise to a symplectic form that provides the commutation relations between the amplitudes of the Goldstone-Gibbs modes, hence behaving as conjugate coordinate-momenta. The dynamics inherits this correspondence, where the amplitudes of the Gibbs modes behave as conserved momenta, since they represent the quantum fluctuations of the corresponding charges, while the amplitudes of the Goldstone modes display a linear time dependence, with a velocity linear in the momenta. Our results for simultaneous symmetry breaking in a stationary state thus generalize the original works of Refs. [55, 56] to an arbitrary number of broken symmetries, and provide a deeper understanding of the nature of the Goldstone-Gibbs modes.
@@
As an example, we apply our formalism to a cnoidal wave in a superfluid, a model that has been used to understand fundamental aspects of supersolidity [57], validating our theoretical results. An important conclusion of this study is that periodic motion along a closed orbit spawned by the broken-symmetry generators may result in the misidentification of time-crystalline behavior, not representing a genuine symmetry-breaking of time-translation invariance.
@@
The generality of the developed framework allows its application for the study of simultaneous symmetry breaking in Floquet states, where the NG modes are now translated into Floquet-Nambu-Goldstone (FNG) modes with zero quasi-energy. In particular, since spontaneous Floquet states break continuous time-translation symmetry, they present a genuine temporal FNG mode. For a complete characterization of the problem, we develop the $(t,\phi)$ formalism, denoted in this way because of its analogy with the standard $(t,t^{\prime})$ formalism [58, 59]. In combination with the generalized Gibbs ensemble, the $(t,\phi)$ formalism leads to a thermodynamic description analogous to that of stationary states, which we term as Floquet thermodynamics. For spontaneous Floquet states, this is a direct consequence of energy conservation, which implies the existence of a conserved Floquet charge. Conventional Floquet states also admit a thermodynamic description in terms of the Floquet enthalpy, the Legendre transform of the energy with respect to the Floquet charge. This is because driven systems are isoperiodic, i.e., they operate at an externally fixed frequency, which is the conjugate variable of the Floquet charge, in analogy with isobaric or isothermal systems where the environment fixes the pressure or the temperature, respectively. The temporal FNG mode and the Floquet charge are thus the distinctive features of a spontaneous Floquet state, absent in driven systems.
@@
The quantization of the Goldstone-Gibbs modes is performed along the same lines as in the stationary case, described by a Berry-Gibbs connection that results in a similar coordinate-momentum correspondence. However, a novel feature is that the amplitude of the temporal FNG mode can be regarded as an effective time coordinate, whose conjugate momentum represents the energy fluctuations. Therefore, the quantum amplitude of the temporal FNG mode represents a unique realization of a time operator in Quantum Mechanics.
@@
Finally, we apply our formalism to the CES state, computing its quantum fluctuations with the help of the Truncated Wigner method [60, 61]. Theory predicts that, at long times, the temporal FNG mode dominates the density-density correlations, observing an excellent agreement with the numerical results. Based on these results, we propose a feasible experimental scheme to observe the temporal FNG mode of the CES state.
@@
The article is arranged as follows. Section 2 discusses how the general results from Goldstone theorem are recovered in variational approaches in many-body systems. Section 3 addresses simultaneous symmetry breaking in stationary states within the generalized Gibbs ensemble, giving rise to the concept of Berry-Gibbs connection. Section 4 applies the formalism of Sec. 3 to a cnoidal wave in a superfluid. Section 5 extends the results of Sec. 2 to Floquet states, both spontaneous and conventional. Section 6 discusses simultaneous symmetry breaking in Floquet systems using the $(t,\phi)$ formalism within the generalized Gibbs ensemble. This leads to the emergence of Floquet thermodynamics, Sec. 6.2, and of a time operator, Sec. 6.4. Section 7 applies the formalism of Sec. 6 to study the quantum fluctuations of the CES state and proposes an experimental setup for the observation of the temporal FNG mode. Technical details are presented in the Appendix.
2 Variational Goldstone Theorem
@@
Originally derived within the framework of relativistic quantum field theory [62], the general idea behind Goldstone theorem is universal, namely, solutions which break some continuous symmetry of the equations of motion include a linear zero mode associated to the broken symmetry. This feature is even present in classical mechanics. Consider a particle of mass $m$ moving within a potential $V(\mathbf{x})$ which is invariant under some continuous symmetry transformation of the coordinates, $V(\mathbf{x}^{\prime})=V(\mathbf{x})$, $\mathbf{x}^{\prime}=\mathbf{x}^{\prime}(\mathbf{x},\alpha)$, where $\alpha$ parameterizes the transformation in such a way that $\alpha=0$ corresponds to the identity transformation, $\mathbf{x}^{\prime}(\mathbf{x},0)=\mathbf{x}$. As standard in Lie theory, the symmetry transformation can be expanded as
@@
where the vector $\boldsymbol{\xi}=-iL\mathbf{x}$ describes the symmetry transformation at the infinitesimal level and $L$ is the generator of the continuous transformation on the coordinates. At the same time, Noether theorem guarantees the existence of a conserved quantity associated to that symmetry.
Time-independent solutions to the equations of motion are given by the equilibrium points $\mathbf{x}_{0}$ which extremize the potential, $\boldsymbol{\nabla}V(\mathbf{x}_{0})=0$. Linearized motion around equilibrium positions is characterized by the eigenvalues $m\Omega^{2}_{i}$ of the Hessian matrix $\partial^{2}_{ij}V(\mathbf{x}_{0})$. In particular, potential minima are stable equilibrium points, as then all $\Omega_{i}$ are real and represent the frequencies of the normal modes of oscillation.
@@
$\partial^{2}_{ij}V(\mathbf{x}_{0})\xi_{0}^{j}=0,\leavevmode\nobreak\ \boldsymbol{\xi}_{0}=-iL\mathbf{x}_{0},$ (2)
@@
where hereafter we use Einstein convention summation unless otherwise stated. Thus, if the symmetry is broken by the equilibrium position, which means that it is not invariant under the symmetry transformation, $L\mathbf{x}_{0}\neq 0$, then there is a zero-frequency normal mode, the Nambu-Goldstone mode. Qualitatively, this zero mode results from the fact that there is no restoring force along the orbit of the equilibrium point under the symmetry transformation because potential energy is conserved there.
@@
A simple example to visualize this behavior is given by a particle in a 2D $y$-independent potential, $V(x,y)=V(x)$, with $V(0)$ a global minimum and $V^{\prime\prime}(0)=m\omega_{x}^{2}$. In that case, if we consider that the particle is initially placed at the equilibrium position $\mathbf{x}=(0,y_{0})$, small displacements along the $x$-axis lead to oscillations with frequency $\omega_{x}$. However, due to the symmetry of the problem, the momentum along the $y$-axis $p_{y}$ is conserved and any small displacement along this direction leads to an unbounded motion with constant velocity $v_{y}=p_{y}/m$, $y(t)=y_{0}+v_{y}t$, since there is no restoring force in this direction.
[…]
$i\hbar\frac{dX^{i}}{dt}=H^{i}_{j}(\mathbf{X})X^{j},$ (4)
@@
where the vector components $X^{i}$ can be orbital wavefunctions (GP and HF equations), wave-function coefficients (Gutzwiller ansatz), or even both (MCTDH method). The matrix $H^{i}_{j}(\mathbf{X})$ usually depends on the vector $\mathbf{X}$ due to many-body interactions, acting as the effective Hamiltonian of the problem. For the present moment, we take the original Hamiltonian $\hat{H}$ as time-independent, so $H^{i}_{j}(\mathbf{X})$ only depends implicitly on time through $\mathbf{X}$.
@@
Now, we assume that there are stationary solutions of the form $X^{i}(t)=X_{0}^{i}e^{-i\epsilon^{i}t/\hbar}$, satisfying the self-consistent eigenvalue equation
$\epsilon^{i}X_{0}^{i}=H^{i}_{j}(\mathbf{X}_{0})X_{0}^{j}.$ (5)
@@
The spectrum of collective modes associated to this stationary solution is obtained by considering small perturbations around $\mathbf{X}_{0}$, $X^{i}(t)=[X_{0}^{i}+\delta X^{i}(t)]e^{-i\epsilon^{i}t/\hbar}$, which yields a linear equation of the general form
$i\hbar\frac{d\delta X^{i}}{dt}=M^{i}_{j}(\mathbf{X}_{0})\delta X^{j},$ (6)
@@
with $M^{i}_{j}(\mathbf{X}_{0})$ some matrix that depends on the background stationary solution. The eigenvalues of $M^{i}_{j}(\mathbf{X}_{0})$ are typically the energies of the collective modes. This is indeed the case of the Bogoliubov-de Gennes (BdG) equations for both the GP equation [64] and the Gutzwiller ansatz [69], and of the Time-Dependent Hartree-Fock Approximation (TDHFA) for the HF equations [70].
@@
In the same fashion as in the classical example, if the stationary solution $\mathbf{X}_{0}$ spontaneously breaks a continuous symmetry of the problem, so $\mathbf{X}_{0}^{\prime}=\mathbf{X}_{0}^{\prime}(\mathbf{X}_{0},\alpha)=\mathbf{X}_{0}-i\alpha L\mathbf{X}_{0}+\ldots$ is also a stationary solution, then the vector $\mathbf{Y}\equiv-iL\mathbf{X}_{0}\neq 0$ is a zero mode of the linearized equations of motion,
$0=M^{i}_{j}(\mathbf{X}_{0})Y^{j}.$ (7)
@@
This is the NG mode associated to the spontaneous symmetry breaking. We note that this derivation also applies to pseudo-NG modes [71, 72], which emerge from extended symmetries in the equations of motion that were not present in the original Hamiltonian. It can be also applied to few-body or open quantum systems governed by similar self-consistent dynamics.
3 Simultaneous symmetry breaking in stationary states